Solve y''+2y'+10y= with initial values y(0)=0, y'(0)=0

Solve

y''+2y'+10y=\delta(t-\pi) with initial values y(0)=0, y'(0)=0
Answer #1

.

L(y) = y(s)

L(y) = sy(s)-y(0)

L(y'')=s^2y(s)-sy(0)-y'(0)

.

.

y''+2y'+10y=\delta \left(t-\pi \right)

take laplace

L[y'']+2L[y']+10L[y]=L[\delta \left(t-\pi \right)]

s^2y(s)-sy(0)-y'(0)+2[ sy(s)-y(0)]+10y(s)= e^{-\pi s}

s^2y(s)-sy(0)-y'(0)+2 sy(s)-2y(0)+10y(s)= e^{-\pi s}

here we have y(0)=0, and y'(0)=0

s^2y(s)-s(0)-0+2 sy(s)-2(0)+10y(s)= e^{-\pi s}

s^2y(s)-0-0+2 sy(s)-0+10y(s)= e^{-\pi s}

s^2y(s)+2 sy(s)+10y(s)= e^{-\pi s}

y(s)(s2 + 2s+ 10) = e-rs

y(s) = \frac{e^{-\pi s}}{ s^2+2s+10 }

apply inverse laplace rule  F (s) ol a=π s2 2s 10 s2 +2s 10

.

now find inverse laplace of

L^{-1}\left\{\frac{1}{s^2+2s+10}\right\}

=L^{-1}\left\{\frac{1}{s^2+2s+1+9}\right\}

=L-1 { (s+1)2 + 9

apply inverse laplace rule    L-1 {F (s-a)) = eat f (t)

=e^{-t}L^{-1}\left\{\frac{1}{s^2+9}\right\}

=e^{-t}L^{-1}\left\{\frac{1}{s^2+3^2}\right\}

=e^{-t } \cdot \frac{1}{3} \cdot L^{-1} \left\{\frac{3}{s^2+3^2}\right\}

=e^{-t } \cdot \frac{1}{3} \cdot \sin \left(3t\right)

= e^{-t}\frac{1}{3}\sin \left(3t\right)

here we have  a=\pi so final answer is

{\color{Red} y=\text{u}\left(t-\pi \right)e^{-\left(t-\pi \right)}\frac{1}{3}\sin \left(3\left(t-\pi \right)\right)}

or

{\color{Red} y=\text{u}_{\pi}\left(t \right)e^{-\left(t-\pi \right)}\frac{1}{3}\sin \left(3\left(t-\pi \right)\right)}

.

where u(t) is Heaviside step function